# Measure-Theoretic Time-Delay Embedding

### v1

@@ 3.2 Differentiable maps on $\mathcal{P}_{2}(M)$
Moreover, a map between two manifolds is differentiable if it is locally differentiable in any chart. This definition cannot be directly translated to $\mathcal{P}_{2}(M)$ since it involves a pointwise evaluation of the differential map in any given chart, whereas the tangent vectors in $\mathcal{P}_{2}(M)$ are only defined almost everywhere. Therefore, we use an equivalent definition of differentiability [30].
@@
Next, we consider the particular situation where the map $F:\mathcal{P}_{2}(M)\rightarrow\mathcal{P}_{2}(N)$ is the pushforward of some $f:M\rightarrow N$. This is exactly the case for the measure-theoretic delay-embedding map $\Psi_{h,\phi_{\tau}}:=\Phi_{h,\phi_{\tau}}\#$. Since the classic delay-embedding map $\Phi_{h,\phi_{\tau}}$ is invertible, we will specifically analyze invertible $f$. A generalization of Theorem 3 below can be found in [31]. For completeness, we provide a full proof of Theorem 3 in the supplementary materials.
@@
Let $F=f\#$ as described above and assume $f:(M,g)\rightarrow(N,q)$ is continuously differentiable, invertible and proper such that $\sup_{x\in M}\|df_{x}\|<\infty$ where $q_{y}$ is the Riemannian metric on $N$. Then $F$ is differentiable (in the sense of Definition 8) and $dF_{\rho}=P^{F(\rho)}\widetilde{dF}_{\rho}$, where
@@ 3.4 Statement of the main theorem
Let $\phi_{t}:M\rightarrow M$ be a dynamical system on a compact $d$-dimensional manifold $M$ such that its vector field satisfies the conditions of Theorem 1. For an observable $h\in C^{2}(M,\mathbb{R})$, define the delay embedding map $\Phi_{h,\phi_{\tau}}:M\rightarrow\mathbb{R}^{m}$ as in (1) and let $\Psi_{h,\phi_{\tau}}:\mathcal{P}_{2}(M)\rightarrow\mathcal{P}_{2}(\mathbb{R}^{m})$ be its pushforward, i.e., $\Psi_{h,\phi_{\tau}}\rho=\Phi_{h,\phi_{\tau}}\#\rho$. Then, if $m\geq 2d+1$, it is a generic property that $\Psi_{h,\phi_{\tau}}$ is an embedding of $\mathcal{P}_{2}(M)$ into $\mathcal{P}_{2}(\mathbb{R}^{m})$ (in the sense of Definition 9).
@@
where $\nu=F({\rho})=f\#\rho$, and $q_{y}$ denotes the Riemannian inner product in $N$. Further derivation shows that
$\int_{N}q_{y}\Big{(}\nabla\varsigma(y),dF_{\rho}(v)(y)-dF_{\rho}(w)(y)\Big{)}{\mathrm{d}}\nu(y)$
$=$ $\int_{f(M)}q_{y}\Big{(}\nabla\varsigma(y),dF_{\rho}(v-w)(y)\Big{)}{\mathrm{d}}\nu(y)$
$=$ $\int_{f(M)}q_{y}\Big{(}\nabla\varsigma(y),\widetilde{dF}_{\rho}(y)(v-w)(y)\Big{)}{\mathrm{d}}\nu(y)$
@@
$=$ $\int_{M}q_{y}\Big{(}\nabla\varsigma(f(x)),df_{x}(v-w)(x)\Big{)}{\mathrm{d}}\rho(x)$
@@
The only claim left to show is that the $\varsigma$ defined above lies in $\mathcal{D}(N)$. To show $\varsigma$ is compactly supported, we find
$\text{supp}(\varsigma)=\overline{\{x:\varsigma(x)\neq 0\}}$ =$=\overline{\{x:(\varphi\circ f^{-1})(x)\neq 0\}}$
$\subseteq\overline{\{x:f^{-1}(x)\in\text{supp}(\varphi)\}}$
$=\overline{f(\text{supp}(\varphi))}.$
Since $f$ is a homeomorphism, it maps compact sets to compact sets. Additionally, a closed subset of a compact set is compact. Hence, we have that $\text{supp}(\varsigma)$ is compact. Moreover, $\varsigma\in\mathcal{C}^{1}(N)$ because $\varphi\in\mathcal{C}^{1}$ and $f\in\mathcal{C}^{1}(M,N)$. We conclude our proof with $\varsigma\in\mathcal{D}(N)$. ∎
@@ 4.1.1 Motivation
We begin by considering samples $\{x_{i}\}$ along a trajectory of a smooth flow $\phi_{t}:M\to M$, where $M\subseteq\mathbb{R}^{n}$ is a smooth compact $d$-dimensional manifold. In applications, samples $\{x_{i}\}$ of the full state can rarely be observed directly, and instead, one may only have access to the values $\{h(x_{i})\}$ of an observable along the trajectory. For suitably chosen time-delay parameters $m\in\mathbb{N}$ and $\tau>0$, the map $\Phi=\Phi_{h,\phi_{\tau}}$ is an embedding of $M$, and one can form the time-delayed trajectory $\{\Phi(x_{i})\}$. It also holds that $\{\Phi(x_{i})\}$ and $\{x_{i}\}$ are related pointwisely by the smooth deterministic map $\Phi^{-1}:\Phi(M)\to\mathbb{R}^{n}$. Thus, if one can learn the reconstruction map $\Phi^{-1}$ from the paired data $\{(x_{i},\Phi(x_{i}))\}$, then the history of the one-dimensional time-series $\{h(x_{i})\}$ can be used to forecast the entire $n$-dimensional trajectory $\{x_{i}\}.$
@@ 4.1.2 The Measure-Theoretic Loss
$\mathcal{L}_{\text{p}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\|x_{i}-\mathcal{R}_{\theta}(\Phi(x_{i}))\|^{2}_{2}\,.$ (13)
@@
where $\mathfrak{D}:\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathcal{P}_{2}(\mathbb{R}^{n})\to\mathbb{R}$ is a metric or divergence on the space of probability measures. Theorem 4 indicates that for a suitable parameterization of $\mathcal{R}_{\theta}$, the loss (14) can indeed be reduced to zero in a noise-free setting. Moreover, while in the pointwise loss ((13)) we seek to recover a map between $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, in (14) we instead search for a map between the corresponding probability spaces $\mathcal{P}_{2}(\mathbb{R}^{m})$ and $\mathcal{P}_{2}(\mathbb{R}^{n}),$ which is parameterized by the pushforward of some function that maps $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$. Theorem 4 guarantees the existence of such a map between the probability spaces $\mathcal{P}_{2}(\mathbb{R}^{m})$ and $\mathcal{P}_{2}(\mathbb{R}^{n})$ with suitable regularity properties, i.e., it is a smooth embedding in the measure-theoretic sense discussed in Section 3.
@@ 4.1.3 Discussion
Enforcing the measure-theoretic objective (14) bears similarities to the approaches in [29, 3], where the reconstruction map $\mathcal{R}_{\theta}$ is learned by averaging the full state over each cluster in the reconstruction space and then linearly interpolating between these averages in the delay coordinate. While these methods ensure that $\mathcal{R}_{\theta}\#(\Phi\#\mu_{i})$ and $\mu_{i}$ agree in expectation, our measure-theoretic approach is designed to match not only the expectation but also all moments of the measures (see (14)).
Here, we discuss the relationship between the pointwise and measure-theoretic loss functions in more detail. If the distributional loss $\mathcal{L}_{\text{m}}$ ((14)) is reduced to zero, then in general, the pointwise loss $\mathcal{L}_{\text{p}}$ ((13)) may still be large. As the diameter of each partition element $C_{i}$ decreases, this discrepancy becomes small, and in the limit when $\mu_{i}=\delta_{x_{i}}$, the loss functions $\mathcal{L}_{\text{p}}$ and $\mathcal{L}_{\text{m}}$ are equivalent for a suitable choice of $\mathfrak{D}$, e.g., the squared Wasserstein distance $\mathfrak{D}=W_{2}^{2}$. Therefore, $\mathcal{L}_{\text{m}}$ should be viewed as a relaxation of $\mathcal{L}_{\text{p}}$, where the diameter of each partition element controls the extent to which pointwise errors in the measure-based reconstruction are permitted. In practice, the partition elements’ diameter should be chosen according to the number of data points and the amount of noise present; see [29, Fig. 6] for a similar discussion.
@@
Throughout our numerical experiments, $\mathcal{R}_{\theta}$ is parameterized as a standard feed forward neural network, and the weights and biases $\theta$ are optimized using Adam [28]. We choose $\mathfrak{D}$ to be the Maximum Mean Discrepancy (MMD) [23] based on either the polynomial kernel $k_{\text{p}}(x,y)=-\|x-y\|_{2}$ or the Gaussian kernel $k_{\text{g}}(x,y)=\exp(-\|x-y\|^{2}_{2}/2\sigma^{2})$. We note that the MMD based upon the polynomial kernel $k_{\text{p}}(x,y)$ is also known as the Energy Distance MMD. We use the Geomloss library to compute $\mathfrak{D}$, which is fully compatible with PyTorch’s autograd engine [19]. We also use teaspoon  [36] to inform our selection of the embedding parameters $\tau>0$ and $m\in\mathbb{N}$ using both the mutual information [20] and Cao’s method [13].
@@ 4.2 First Examples: Noisy Chaotic Attractors
We begin by studying our measure-theoretic approach to state reconstruction on the Lorenz-63 system [52], the Rössler system [43], and a four-dimensional Lotka–Volterra model [53]. For these dynamical systems, we select standard values for the systems’ parameters, which are known to produce chaotic trajectories; see the supplementary details for the system equations and precise parameter choices.
The task is to reconstruct the full state of these systems using partial observations. For each system, we first simulate a long trajectory, form the delayed state based on a scalar observable, and split the data into training and testing components. As explained in Section 4.1, the measures $\{\Phi\#\mu_{i}\}_{i=1}^{K}$ are given by conditioning a long trajectory in time-delay coordinates on various regions of the attractor, which in practice is implemented by a k-means clustering algorithm.
We remark that both the pointwise approach (13) and the measure-based approach (14) can achieve accurate reconstructions when the data is noise-free; see the supplementary materials for an experiment demonstrating the measure-based approach applied to clean data. However, our method proves particularly advantageous when dealing with sparse and noisy data. To demonstrate this, we compare the performance of the measure-based method with pointwise matching in Fig. 2 using imperfect data. In these tests, extrinsic noise is applied to the entire state, including the time series that forms the time-delay coordinates. From these corrupted inputs and outputs, we learn the full-state reconstruction map. Although neither method is expected to achieve perfect reconstruction, the measure-based approach yields smoother results, whereas the pointwise approach tends to overfit the noise.

### v2

@@ 3.2 Differentiable Maps on $\mathcal{P}_{2}(M)$
Moreover, a map between two manifolds is differentiable if it is locally differentiable in any chart. This definition cannot be directly translated to $\mathcal{P}_{2}(M)$ since it involves a pointwise evaluation of the differential map in any given chart, whereas the tangent vectors in $\mathcal{P}_{2}(M)$ are only defined almost everywhere. Therefore, we use an equivalent definition of differentiability [47, Chapter 3].
@@
Next, we consider the particular situation where the map $F:\mathcal{P}_{2}(M)\rightarrow\mathcal{P}_{2}(N)$ is the pushforward of some $f:M\rightarrow N$. This is exactly the case for the measure-theoretic delay-embedding map $\Psi_{h,\phi_{\tau}}:=\Phi_{h,\phi_{\tau}}\#$. Since the classic delay-embedding map $\Phi_{h,\phi_{\tau}}$ is invertible, we will specifically analyze invertible $f$. A generalization of Theorem 7 below can be found in [48]. For completeness, we provide a full proof of Theorem 7 in Appendix A.
@@
Let F=f#F=f\# as described above and assume $f:(M,g)\rightarrow(N,q)$ is continuously differentiable, invertible and proper such that $\sup_{x\in M}\|df_{x}\|<\infty$. Then $F$ is differentiable (in the sense of Definition 8) and $dF_{\rho}=P^{F(\rho)}\widetilde{dF}_{\rho}$, where
@@ 3.4 Statement of the Main Theorem
Let $\phi_{t}:M\rightarrow M$ be a dynamical system on a compact $d$-dimensional manifold $M$ such that its vector field satisfies the conditions of Theorem 1. For an observable $h\in\mathcal{C}^{2}(M,\mathbb{R})$, define the delay embedding map $\Phi_{h,\phi_{\tau}}:M\rightarrow\mathbb{R}^{m}$ as in (1) and let $\Psi_{h,\phi_{\tau}}:\mathcal{P}_{2}(M)\rightarrow\mathcal{P}_{2}(\mathbb{R}^{m})$ be its pushforward, i.e., $\Psi_{h,\phi_{\tau}}\rho=\Phi_{h,\phi_{\tau}}\#\rho$. Then, if $m\geq 2d+1$, it is a generic property that $\Psi_{h,\phi_{\tau}}$ is an embedding of $\mathcal{P}_{2}(M)$ into $\mathcal{P}_{2}(\mathbb{R}^{m})$ (in the sense of Definition 9).
@@
where $\nu=F({\rho})=f\#\rho$, and $q_{y}$ denotes the Riemannian inner product in $N$ at the point $y\in N$. Further derivation shows that
$\int_{N}q_{y}\Big(\nabla\varsigma(y),dF_{\rho}(v)(y)-dF_{\rho}(w)(y)\Big){\mathrm{d}}\nu(y)$ $\text{ By the linearity of }dF_{\rho}$
$=$ $\int_{f(M)}q_{y}\Big(\nabla\varsigma(y),dF_{\rho}(v-w)(y)\Big){\mathrm{d}}\nu(y)$ $\text{ Since }dF_{\rho}-\widetilde{dF}_{\rho}\text{ is divergence free}$
$=$ $\int_{f(M)}q_{y}\Big(\nabla\varsigma(y),\widetilde{dF}_{\rho}(v-w)(y)\Big){\mathrm{d}}\nu(y)$ Using Theorem 7
@@
$=$ $\int_{M}q_{f(x)}\Big(\nabla\varsigma(f(x)),df_{x}(v-w)(x)\Big){\mathrm{d}}\rho(x)$ Using Proposition 4
@@
The only claim left to show is that the $\varsigma$ defined above lies in $\mathcal{D}(N)$. To show $\varsigma$ is compactly supported, note that
$\{x:\varsigma(x)\neq 0\}$ $=\{x:\varphi(f^{-1}(x))\neq 0\}$
$=\{x:f^{-1}(x)\in\{y:\varphi(y)\neq 0\}\}$
$=f\bigl(\{y:\varphi(y)\neq 0\}\bigr).$
Taking closures and using the fact that a homeomorphism preserves closures, we get $\text{supp}(\varsigma)=f(\text{supp}(\varphi))$. Also, since $f$ is a homeomorphism, it maps compact sets to compact sets, so $\text{supp}(\varsigma)$ is compact. Moreover, $\varsigma\in\mathcal{C}^{1}(N)$ because $\varphi\in\mathcal{C}^{1}$ and $f^{-1}\in\mathcal{C}^{1}(M,N)$. We conclude our proof with $\varsigma\in\mathcal{D}(N)$. ∎
@@ 4.1.1 Motivation
We begin by considering samples $\{x_{i}\}$ along a trajectory of a smooth flow $\phi_{t}:M\to M$, where $M\subseteq\mathbb{R}^{n}$ is a smooth compact $d$-dimensional manifold. In applications, samples $\{x_{i}\}$ of the full state can rarely be observed directly, and instead, one may only have access to the values $\{h(x_{i})\}$ of an observable along the trajectory. For suitably chosen time-delay parameters $m\in\mathbb{N}$ and $\tau>0$, the map $\Phi=\Phi_{h,\phi_{\tau}}$ is an embedding of $M$, and one can form the time-delayed trajectory $\{\Phi(x_{i})\}$. It also holds that $\{\Phi(x_{i})\}$ and $\{x_{i}\}$ are related pointwisely by the smooth deterministic map $\Phi^{-1}:\Phi(M)\to\mathbb{R}^{n}$. Thus, if one can learn the reconstruction map $\Phi^{-1}$ from the paired data $\{(x_{i},\Phi(x_{i}))\}$, then the history of the one-dimensional time-series $\{h(x_{i})\}$ can be used to reconstruct the entire $n$-dimensional trajectory $\{x_{i}\}.$
This approach is useful in applications where only limited training data from the full state is available [49]. In particular, it may be possible to access the full state of a high-dimensional system for a short period of time, but it may be prohibitively expensive or challenging to continue measuring it directly at later times. In such a setup, one may build a surrogate mapping from delay embeddings of a partially observed state variable to the full state. Thus, expensive measurements of the full state need not be recorded, and one can instead reconstruct the full state from easier-to-access partial observations. This problem is common in the surrogate modeling of fluid flows where one builds a full-state reconstruction map using only a few sparsely distributed sensors [50, 51].
In certain applications, no data from the full state can be observed. In these cases, one can still predict the evolution of the one-dimensional time series $\{h(x_{i})\}$ by learning the map
$\mathcal{T}:\Phi(M)\to\Phi(M),\qquad\mathcal{T}(\Phi(x_{i}))=\Phi(x_{i+1}),$ (13)
where here we have assumed that the samples $\{x_{i}\}$ are obtained uniformly in time. Then, the delay state can be predicted by iterated composition with $\mathcal{T}$, and the forecasted time series can be recovered by projecting to a single axis.
@@ 4.1.2 The Measure-Theoretic Loss
$\mathcal{L}_{\text{p}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}|x_{i}-\mathcal{R}_{\theta}(\Phi(x_{i}))|^{2}\,.$ (14)
@@
where $\mathfrak{D}:\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathcal{P}_{2}(\mathbb{R}^{n})\to\mathbb{R}$ is a metric or divergence on the space of probability measures. Theorem 8 indicates that for a suitable parameterization of $\mathcal{R}_{\theta}$, the loss (15) can indeed be reduced to zero in a noise-free setting. Moreover, while in the pointwise loss (14) we seek to recover a map between $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, in (15) we instead search for a map between the corresponding probability spaces $\mathcal{P}_{2}(\mathbb{R}^{m})$ and $\mathcal{P}_{2}(\mathbb{R}^{n}),$ which is parameterized by the pushforward of some function that maps $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$. Theorem 8 guarantees the existence of such a map between the probability spaces $\mathcal{P}_{2}(\mathbb{R}^{m})$ and $\mathcal{P}_{2}(\mathbb{R}^{n})$ with suitable regularity properties, i.e., it is a smooth embedding in the measure-theoretic sense discussed in Section 3.
@@
We also remark that the loss functions (14) and (15) can be modified to account for the case when one is interested in learning the dynamics $\Phi(x_{i})\mapsto\Phi(x_{i+1})$ in delay coordinates. For this related problem, the pointwise loss becomes
$\mathcal{L}_{\textrm{p}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}|\Phi(x_{i+1})-\mathcal{T}_{\theta}(\Phi(x_{i}))|^{2},$ (17)
where $\mathcal{T}_{\theta}$ is a neural-network parameterized flow map representing the dynamics in time-delay coordinates. Similarly, the measure-theoretic analog is
$\mathcal{L}_{\text{m}}(\theta)=\frac{1}{K}\sum_{i=1}^{K}\mathfrak{D}\left(\mathcal{T}\#(\Phi\#\mu_{i}),\mathcal{T}_{\theta}\#(\Phi\#\mu_{i})\right),$ (18)
where $\mathcal{T}$ is defined in (13). Notably, both (17) and (18) can be evaluated without having direct access to $\mathcal{T}$ and without observing the full state of the dynamical system. While not explored in this work, one can also leverage the inherent shift-based structure of the delay-coordinate dynamics $\mathcal{T}$ to restrict the hypothesis space of flow maps in learning $\mathcal{T}_{\theta}$.
@@ 4.1.3 Discussion
Enforcing the measure-theoretic objective (15) bears similarities to the approaches in [13, 53], where the reconstruction map $\mathcal{R}_{\theta}$ is learned by averaging the full state over each cluster in the reconstruction space and then linearly interpolating between these averages in the delay coordinate. While these methods ensure that $\mathcal{R}_{\theta}\#(\Phi\#\mu_{i})$ and $\mu_{i}$ agree in expectation, our measure-theoretic approach is designed to match not only the expectation but also all moments of the measures (see Eqn. (15)).
Here, we discuss the relationship between the pointwise and measure-theoretic loss functions in more detail. If the distributional loss $\mathcal{L}_{\text{m}}$ (see Eqn. (15)) is reduced to zero, then in general, the pointwise loss $\mathcal{L}_{\text{p}}$ (see Eqn. (14)) may still be large. As the diameter of each partition element $C_{i}$ decreases, this discrepancy becomes small, and in the limit when $\mu_{i}=\delta_{x_{i}}$, the loss functions $\mathcal{L}_{\text{p}}$ and $\mathcal{L}_{\text{m}}$ are equivalent for a suitable choice of $\mathfrak{D}$, e.g., the squared Wasserstein distance $\mathfrak{D}=W_{2}^{2}$. Therefore, $\mathcal{L}_{\text{m}}$ should be viewed as a relaxation of $\mathcal{L}_{\text{p}}$, where the diameter of each partition element controls the extent to which pointwise errors in the measure-based reconstruction are permitted. In practice, the partition elements’ diameter should be chosen according to the number of data points and the amount of noise present; see [13, Fig. 6] for a similar discussion.
@@
Both the diameter of the cells $C_{i}$ and the number of samples $N_{i}$ residing in each cell contribute potential sources of error when using the measure-based loss (15). As the number of samples $N_{i}$ increases and the diameters of the cells $C_{i}$ decrease, the accuracy of the measure-based reconstruction is expected to increase in an ideal noise-free setting. When the available data is noise-free, the use of empirical measures supported on finite diameter cells may unnecessarily hinder accuracy and the pointwise loss (14) is generally preferred. However, when the data is corrupted by noise, the pointwise loss (14) tends to overfit the observed data, while we find the flexibility of the measure-based matching (15) to be more robust. We also remark that the construction (16) of empirical measures is only one possible way of forming measure-valued data $\mu_{i}$ for training the loss (15). In this work, we choose to consider localized measures supported on Voronoi cells, as the cell diameters can be tuned to effectively balance approximation accuracy and noise-tolerance.
Throughout our numerical experiments, $\mathcal{R}_{\theta}$ is parameterized as a standard feed forward neural network, and the weights and biases $\theta$ are optimized using Adam [57]. Similarly, the delay-coordinate dynamics $\mathcal{T}_{\theta}$ are defined by the flow map of a neural network-parameterized vector field. We choose $\mathfrak{D}$ to be the Maximum Mean Discrepancy (MMD) [58] based on either the kernel $k_{1}(x,y)=-\|x-y\|_{2}$ or the Gaussian kernel $k_{2}(x,y)=\exp(-\|x-y\|^{2}_{2}/2\sigma^{2})$. We note that the MMD based upon the kernel $k_{1}(x,y)$ is also known as the Energy Distance MMD. We use the Geomloss library to compute $\mathfrak{D}$, which is fully compatible with PyTorch’s autograd engine [59]. There are several times in our numerical experiments where the chosen embedding dimension $m\in\mathbb{N}$ does not satisfy the inequality $m\geq 2d+1$ introduced in Theorem 1. While Theorem 1 provides an upper bound on the minimum required embedding dimension, it is often the case that lower-dimensional embeddings can be achieved. Several works have studied the problem of estimating such an embedding dimension numerically from partially observed time-series data. In this work, we use teaspoon [60] to inform our selection of the embedding parameters $\tau>0$ and $m\in\mathbb{N}$ using both the mutual information [41] and Cao’s method [40]. Notably, [40] estimates the minimum embedding dimension of the Lorenz-63 system to be $m=3$, while the upper bound in Theorem 1 suggests taking $m\geq 7$.
@@ 4.2 First Examples: Noisy Chaotic Attractors
4.2.1 State Reconstruction
We begin by studying our measure-theoretic approach to state reconstruction on the Lorenz-63 system [61], the Rössler system [62], and a four-dimensional Lotka–Volterra model [63]. For these dynamical systems, we select standard values for the systems’ parameters, which are known to produce chaotic trajectories; see Appendix B for the system equations and precise parameter choices.
The task is to reconstruct the full state of these systems using partial observations. In particular, we aim to learn the inverse embedding map $\Phi^{-1}$ using input-output training data from both the delay state of a partial observation and the system’s full state. For each system, we first simulate a long trajectory, form the delayed state based on a scalar observable, and split the data into training and testing components. As explained in Section 4.1, the measures $\{\Phi\#\mu_{i}\}_{i=1}^{K}$ are given by conditioning a long trajectory in time-delay coordinates on various regions of the attractor, which in practice is implemented by a k-means clustering algorithm.
We remark that both the pointwise approach (14) and the measure-based approach (15) can achieve accurate reconstructions when the data is noise-free; see Appendix B for an experiment demonstrating the measure-based approach applied to clean data. However, our method proves particularly advantageous when dealing with sparse and noisy data. To demonstrate this, we compare the performance of the measure-based method with pointwise matching in Fig. 2 using imperfect data. In these tests, extrinsic noise is applied to the entire state, including the time series that forms the time-delay coordinates. From these corrupted inputs and outputs, we learn the full-state reconstruction map. Although neither method is expected to achieve perfect reconstruction, the measure-based approach yields smoother results, whereas the pointwise approach tends to overfit the noise.